OpenStudy (anonymous):

PLEASEEE HELP ME :( 1.)To open these doors, you must speak the function in standard form... *One function, h(x), with two complex solutions. ....... i need an example of this im confused :(

OpenStudy (jdoe0001):

hmm have you covered the quadratic formula?

OpenStudy (jdoe0001):

... or do you know what a "complex number" is, and have you done quadratic equations yet?

OpenStudy (anonymous):

yes i know that a complex number is a combination of a real and imaginary number, and yes.

OpenStudy (jdoe0001):

ok... so you just need to make a quadratic equation with 2 complex solutions so, the way to do it, will be to first, pick your 2 complex solutions, so which one do you think we can use?

OpenStudy (jdoe0001):

got any 2 complex expressions we can use, pick any 2

OpenStudy (anonymous):

sorry my computer froze

OpenStudy (anonymous):

but is -6i and -12i???? work

OpenStudy (jdoe0001):

sure

OpenStudy (anonymous):

okay so what is the next step

OpenStudy (jdoe0001):

ohh... one sec... yes so we know that the solutions are x = -6i x = -12i so \(\bf x=-6i\implies x+6i=0\implies (x+6i)=0\\ \quad \\ x=-12i\implies x+12i=0\implies (x+12i)=0\\ \quad \\ \textit{so our polynomial will be } (x+6i)(x+12i)=0\)

OpenStudy (anonymous):

i would not choose \(6i\) and \(12i\) because you will not get a polynomial with real coefficients it is much easier than that

OpenStudy (jdoe0001):

hmm

OpenStudy (anonymous):

since \(x^2\geq 0\) for any \(x\) you can pick something like \[x^2+4\] because \(x^2+4\geq 4\) and so has no zeros

OpenStudy (jdoe0001):

ohhh... right...

OpenStudy (anonymous):

so can i do x^2+4 and x^2+8

OpenStudy (anonymous):

thank you all

OpenStudy (anonymous):

yes you can

OpenStudy (jdoe0001):

anyhow vickysecret , my bad, the pair usually will be a conjugate version of complex so let's use say.... -6i and its conjugate +6i instead \(\bf x=-6i\implies x+6i=0\implies (x+6i)=0\\ \quad \\ x=6i\implies x-6i=0\implies (x-6i)=0\\ \quad \\ \textit{so our polynomial will be } (x+6i)(x-6i)=0\\ \quad \\ \textit{recall that }(a-b)(a+b) = a^2-b^2\qquad thus\\ \quad \\ (x+6i)(x-6i)=0\implies x^2-(6i)^2=0\implies x^2-6^2(i^2)=0\\ \quad \\ x^2-36(-1)=0\implies x^2+36=0\)

OpenStudy (anonymous):

okay that was alot, so is the official answer x^2+36=0

OpenStudy (jdoe0001):

if we use -6i and +6i, yes what you do is, from the solution, you get a polynomial, then multiply both polynomials to get the original function whose solutions are those

OpenStudy (anonymous):

okay i understand now, thank you so much

OpenStudy (jdoe0001):

yw